# Stochastic gravitational wave from graviton bremsstrahlung in inflaton decay into massive spin 3/2 particles

### v1

@@ I Introduction
In the present paper we discuss graviton bremsstrahlung in the decay of inflaton into massive spin 3/2 particles. Spin 3/2 particles are one of the first dark matter candidates to be ever considered. Though the spin 3/2 particles appear naturally in local supersymmetry, or supergravity theories, as early as 1941 Rarita and Schwinger proposed a Lagrangian of a relativistic free spin 3/2 field Rarita:1941mf based on the earlier works by Fierz and Pauli Fierz:1939ix . The spin 3/2, often known as Rarita-Schwinger (RS) field, is a vector-valued spinor field and is denoted as $\psi_{\mu}$ where $\mu$ is the Lorentz index and the spinor index is usually suppressed. Recently, the gravitational production of spin 3/2 particles has been discussed in Refs. Kaneta:2023uwi ; Casagrande:2023fjk ; Kolb:2023ydq ; Garcia:2020hyo .
@@ V Reheating
After inflation ends, the decay of inflaton to the pair of RS particles and gravitons reheat the universe and produces a stochastic GW background. The evolutions of inflaton’s energy density, $\rho_{\phi}$, and the radiation energy density, $\rho_{R}$, are given by a set of Boltzmann equations
@@
The free parameters of our Lagrangian are the couplings $\lambda_{s,p}$ and the RS particle mass $m_{3/2}$. As the operators $\bar{\psi}_{\mu}\psi^{\mu}\phi$ and $\bar{\psi}_{\mu}\gamma\psi^{\mu}\phi$ do not interfere in the decay rates, for the rest of our analysis we consider only one of the couplings to be non-zero. Also, for the decays to take place $m_{3/2}<m_{\phi}/2$.
@@
The evolutions of $\rho_{\phi}$ and $\rho_{R}$ as a function of the scale factor is shown in figure 2 for $k=2,4,6,8$. The figure demonstrates that as the energy density $\rho_{\phi}$ is diluted due to dissipation of energy to the daughter field, the $\rho_{R}$ increases and reheats the universe. The intersection of $\rho_{\phi}$ and $\rho_{R}$ is defined as reheating. Depending on the shape of the potential $k$, reheating is achieved at different times.
@@ VI Stochastic GW
Finally, in figure 6 we show the present day GW spectra for different $k$ for two sets of parameters: $\lambda_{p}=10^{-4},\lambda_{s}=0,m_{3/2}=0.05m_{\phi}$ (top) and $\lambda_{p}=10^{-6},\lambda_{s}=0,m_{3/2}=0.05m_{\phi}$ (bottom). The spectra are also compared with the sensitivity curves of some of the future GW detectors that include the ultimate Deci-Hertz Interferometer Gravitational Wave Observatory (uDECIGO) Seto:2001qf ; Kudoh:2005as , the Big Bang Observer (BBO) Crowder:2005nr ; Corbin:2005ny , the Resonance Cavities Berlin:2021txa ; Herman:2022fau , and the Einstein Telescope ET:2019dnz ; Sathyaprakash:2012jk ; Hild:2010id ; Punturo:2010zz . Clearly, the GW signals are out of reach of the future detectors. However, if detector sensitivity improves, the signals may be accessible and they can provide insights into the inflaton potential as well as the parameters of the RS Lagrangian.

### v2

@@ I Introduction
In the present paper we discuss graviton bremsstrahlung in the decay of inflaton into massive spin 3/2 particles. Spin 3/2 particles are one of the first dark matter candidates to be ever considered. Though the spin 3/2 particles appear naturally in local supersymmetry, or supergravity theories, as early as 1941 Rarita and Schwinger proposed a Lagrangian of a relativistic free spin 3/2 field Rarita:1941mf based on the earlier works by Fierz and Pauli Fierz:1939ix . The spin 3/2, often known as Rarita-Schwinger (RS) field, is a vector-valued spinor field and is denoted as $\psi_{\mu}$ where $\mu$ is the Lorentz index and the spinor index is usually suppressed. Recently, the gravitational production of spin 3/2 particles has been discussed in Refs. Kaneta:2023uwi ; Casagrande:2023fjk ; Kolb:2023ydq ; Garcia:2020hyo . In supergravity, inflaton decay into spin-$3/2$ particles can arise through effective interactions determined by the Kähler potential and supersymmetry-breaking sector Endo:2006zj Endo:2007sz .
@@ V Reheating
Although the inflaton decay products considered here are spin-$3/2$ fermions, we follow Refs. Nakayama:2018ptw ; Barman:2023ymn ; Barman:2023rpg ; Kanemura:2023pnv ; Bernal:2023wus ; Tokareva:2023mrt and adopt a phenomenological reheating description in which the energy transferred from inflaton decay is treated as contributing to an effective Standard Model radiation bath. The description allows one to meaningfully speak of a temperature $T$ and an effective number of relativistic degrees of freedom $g_{*}(T)$. At the level of effective field theory, a massive spin-$3/2$ field can admit higher-dimensional non-gravitational couplings to light degrees of freedom Garcia:2020hyo ; Gherghetta:2024tob , unless forbidden by an exact symmetry. Such couplings provide a plausible channel for energy transfer beyond purely gravitational interactions. A particularly favorable setup is non-gravitational coupling of the RS field with the Standard Model through a $d=5$ dimensional operator, $\mathcal{L}_{\rm int}\supset(c_{d}/\Lambda^{d-4})\mathcal{O}_{d}(\psi_{\mu},\rm SM)$ where $c_{d}$ is the Wilson coefficient and $\Lambda$ is the ultraviolet cutoff scale. The resulting decay rate scales as $\Gamma_{3/2}\sim{c_{5}^{2}m_{3/2}^{3}/\Lambda^{2}}$. Efficient thermalization requires the RS field to decay faster than the inflaton inverse-decay rate $\Gamma_{3/2}\gg\Gamma_{\phi}\sim\lambda_{s,p}^{2}\,m_{\phi}$. At the same time, consistency of the effective field theory demands $E/\Lambda\ll 1$, with the characteristic energy scale set by $E\sim m_{\phi}$. Both conditions are simultaneously satisfied in our benchmark scenario. For example, choosing $m_{\phi}=10^{13}\,\mathrm{GeV},m_{3/2}=0.05\,m_{\phi},\lambda_{s,p}=10^{-6},\Lambda=10^{15}\,\mathrm{GeV},c_{5}=1$ one finds $\Gamma_{3/2}\gg\Gamma_{\phi}$ and $m_{\phi}/\Lambda\ll 1$. Therefore, efficient thermalization is possible while remaining well within the regime of validity of the effective field theory. We do not further attempt a microscopic description of the equilibration processes leading to this bath. Instead, we focus on the resulting gravitational wave signal within the standard reheating framework.
In this approach, the post-inflationary dynamics are described by coupled Boltzmann equations for the inflaton, radiation, and gravitational wave energy densities, providing a phenomenological mapping between inflaton decay and the subsequent radiation-dominated epoch. The results presented below should therefore be interpreted within this standard phenomenological reheating framework.
The evolutions of inflaton’s energy density, $\rho_{\phi}$, and the SM radiation energy density, $\rho_{R}$, are given by a set of Boltzmann equations
@@
The free parameters of our Lagrangian are the couplings $\lambda_{s,p}$ and the RS particle mass $m_{3/2}$. As the operators $\bar{\psi}_{\mu}\psi^{\mu}\phi$ and $\bar{\psi}_{\mu}\gamma\psi^{\mu}\phi$ do not interfere in the decay rates, for the rest of our analysis we consider only one of the couplings to be non-zero.
@@
The evolutions of $\rho_{\phi}$ and $\rho_{R}$ as a function of the scale factor is shown in figure 2 for $k=2,4,6,8$. The figure demonstrates that as the energy density $\rho_{\phi}$ is diluted due to dissipation of energy to the daughter field, the $\rho_{R}$ increases and reheats the universe. The intersection of $\rho_{\phi}$ and $\rho_{R}$ is defined as reheating. Depending on the shape of the potential $k$, reheating is achieved at different times. For the decays to take place $m_{3/2}<m_{\phi}/2$. In our analysis, the benchmark values of RS particles is restricted to $m_{3/2}\leq 0.05m_{\phi}$ ensuring that the spin-3/2 fermions are produced in the deep-relativistic to ultrarelativistic regime.
@@ VI Stochastic GW
Finally, in figure 6 we show the present day GW spectra for different $k$ for two sets of parameters: $\lambda_{p}=10^{-4},\lambda_{s}=0,m_{3/2}=0.05m_{\phi}$ (top) and $\lambda_{p}=10^{-6},\lambda_{s}=0,m_{3/2}=0.05m_{\phi}$ (bottom). The spectra are also compared with the sensitivity curves of some of the future GW detectors that include the ultimate Deci-Hertz Interferometer Gravitational Wave Observatory (uDECIGO) Seto:2001qf ; Kudoh:2005as , the Big Bang Observer (BBO) Crowder:2005nr ; Corbin:2005ny , the Resonance Cavities Berlin:2021txa ; Herman:2022fau ; Capdevilla:2024cby , and the Einstein Telescope ET:2019dnz ; Sathyaprakash:2012jk ; Hild:2010id ; Punturo:2010zz . Clearly, the GW signals are out of reach of the future detectors. However, if detector sensitivity improves, the signals may be accessible and they can provide insights into the inflaton potential as well as the parameters of the RS Lagrangian.
@@ Acknowledgment
Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.